Reports 1-1 of 1 Clear search Modify search
terrence.tsang - 18:20 Wednesday 24 July 2019 (9629) Print this report
SR3 IM H OSEM Noise issue temporary fix

With Fabian,

We couldn't figure out what is the source of the presistent noise appearing in the ADC CH9 and CH10 which happen to be used for IM H2 and IM H3. To temporarily fix this, we plugged the IM H OSEM cable to ADC CH28-CH31. So, CH28 is measuring IM H1, CH29 is measuring IM H2 and CH30 is measuring IM H3 while CH31 remains floated unless the 4th channel of the satellite box is connected. We measured the noise level again. Preliminary results show a ~0.02 µm pk-pk noise level in IM H2 which is 10 times better than before. A thorough report will be given tomorrow if theres time for additional measurements. 

After fixing the OSEMs, we should revisit the IM damping filters to properly damp the 0.14 Hz yaw mode.

Search Help